Differentiate the following without solving for y in terms of x.

  1. x2/9 + y2/4 =1
  2. x2/4 + y2=1
  3. x2+ y2/9 =1
1
Expert's answer
2022-04-12T05:51:38-0400
  1. "{\frac {2x} 9}+{\frac {2y} 4}*y'=0\implies {\frac {2x} 9}+{\frac {yy'} 2}=0"
  2. "{\frac {2x} 4}+ 2y*y'=0\implies {\frac x 2}+2yy'=0"
  3. "2x+{\frac {2y} 9}*y'=0"